successfull run record

Yves-Alain Peter yapeter at
Sat Oct 19 16:01:44 PDT 2002

As per your request, please, find below details of today's successfull run:

SiH4: 14
PH3: 86
O2: 115
Temp: 400
Pressure: 326
thickness: around 2000A


Ankur Jain wrote:

> Hello everyone!
>  I have a request to make to labmembers who will be using tylanbpsg in the
> next few days. It would be nice if they could post to this list the
> pressures they get in their runs.
> Thanks!
> Ankur.
>  On Fri, 18 Oct 2002, Ted Berg wrote:
> > hello all, we checked out the system this morning and found the pump
> > screen clogged again. it was changed and tested at process flows:
> > o2-115, siH4-100 pressure at 297 which is normal.  please advise if you
> > have any further problems.  also be aware that thickness is at 25+
> > microns clean is scheduled for 30 so probably next week.ted
> >
-------------- next part --------------
A non-text attachment was scrubbed...
Name: yapeter.vcf
Type: text/x-vcard
Size: 397 bytes
Desc: Card for Yves-Alain Peter
URL: <>

More information about the tylanbpsg mailing list